OpenStudy (anonymous):

A $200.00 sweater has been marked down to $120.00. What percentage discount from the original price does the markdown represent?

OpenStudy (anonymous):

60%

OpenStudy (anonymous):

120/200=0.60 *100%=60%

OpenStudy (anonymous):

New price: $120 Old price: $200 Discount: $80 Discount percentage: $80/$200 = 40%

OpenStudy (anonymous):

Since its 60% then you have 100-60=40%